# Generalizations of the Squircle-Lemniscate Relation and Keplerian Dynamics

### v3

@@ 3. The Lamé-Sinusoidal Spiral Relation
The integral $A=\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ represents the area of the portion of the Lamé curve $x^{2n}+y^{2n}=1$ contained within the first quadrant. An alternative way of computing this area is to use the following parametrization of the Lamé curve:
@@
[Note that these are the same as the polar equations of the sinusoidal spiral restricted to the interval $\left[0,\frac{\pi}{2n}\right]$.] Thus we have
$A$ $=\frac{1}{2}\int_{0}^{\frac{\pi}{2n}}\left(x\frac{dy}{dt}-y\frac{dx}{dt}\right)\,dt$
@@
$A$ $=\frac{1}{2^{1+\frac{1}{n}}}\int_{0}^{\frac{\pi}{n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du$
@@
(11) $A=\frac{1}{2^{1+\frac{1}{n}}}\left[2\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du\right]=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du.$
@@
(12) $A=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}$
@@
$\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}=2^{\frac{1}{n}}A=2^{\frac{1}{n}}\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx.$
@@
$\mathcal{A}\left(x^{2n}+y^{2n}=1\right)$ $\quad=\quad 4\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$
@@
$\mathcal{A}\left(\left(\frac{x}{a}\right)^{2n}+\left(\frac{y}{b}\right)^{2n}=1\right)$ $\quad=\quad 4\int_{0}^{a}b\sqrt[2n]{1-\left(\frac{x}{a}\right)^{2n}}\,dx$

### v4

@@ 3. The Lamé-Sinusoidal Spiral Relation
The integral $\mathcal{A}=\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ represents the area of the portion of the Lamé curve $x^{2n}+y^{2n}=1$ contained within the first quadrant. An alternative way of computing this area is to use the following parametrization of the Lamé curve:
@@
[Note that these are the same as the polar equations of the sinusoidal spiral restricted to the interval $\left[0,\frac{\pi}{2n}\right]$.] Thus by Green’s theorem, we have
$\mathcal{A}$ $=\frac{1}{2}\int_{\partial\{(x,y)\,|\,x^{2n}+y^{2n}\leq 1,\,x\geq 0,\,y\geq 0\}}(x\,dy-y\,dx)$
$=0+0+\frac{1}{2}\int_{0}^{\frac{\pi}{2n}}\left(x\frac{dy}{dt}-y\frac{dx}{dt}\right)\,dt$
@@
$\mathcal{A}$ $=\frac{1}{2^{1+\frac{1}{n}}}\int_{0}^{\frac{\pi}{n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du$
@@
(11) $\mathcal{A}=\frac{1}{2^{1+\frac{1}{n}}}\left[2\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du\right]=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du.$
@@
(12) $\mathcal{A}=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}$
@@
$\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}=2^{\frac{1}{n}}\mathcal{A}=2^{\frac{1}{n}}\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx.$
@@
$\mathcal{A}\left(x^{2n}+y^{2n}\leq 1\right)$ $\quad=\quad 4\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$
@@
$\mathcal{A}\left(\left(\frac{x}{a}\right)^{2n}+\left(\frac{y}{b}\right)^{2n}\leq 1\right)$ $\quad=\quad 4\int_{0}^{a}b\sqrt[2n]{1-\left(\frac{x}{a}\right)^{2n}}\,dx$